# Excitable Nonlinear Opinion Dynamics (E-NOD) for Agile Decision-Making

### v2

@@ II-C Limitation on Tuning of NOD
($K_{u}$ determines region of multi-stability): Let $b=0$ and $(u_{0}^{\dagger},z^{\dagger})$ be either of the two saddle-node bifurcation points of the subcritical pitchfork of NOD (2) for $K_{u}\!>\!\frac{d^{3}}{3\alpha}$. Then $u_{0}^{\dagger}$ is a monotonically decreasing function of $K_{u}$, i.e., $\frac{\partial u_{0}^{\dagger}}{\partial K_{u}}\!<\!0$.
@@
Let $K_{u}^{\dagger}\!>\!\frac{d^{3}}{3\alpha}$ and $f(z,K_{u},u_{0})\!:=\!-d\,z\!+\!\tanh(\alpha z(u_{0}\!+\!K_{u}z^{2}))$. By hypothesis, $f(z^{\dagger},K_{u}^{\dagger},u_{0}^{\dagger})\!=\!0$. We have $\frac{\partial f}{\partial u_{0}}(z^{\dagger},K_{u}^{\dagger},u_{0}^{\dagger})=\alpha z^{\dagger}\tanh^{\prime}(\alpha z^{\dagger}(u_{0}^{\dagger}\!+\!K_{u}^{\dagger}(z^{\dagger})^{2}))\!\neq\!0$, since $z^{\dagger}\!\neq\!0$. Following [19], we use the implicit function theorem to show the existence of $g\!:\!\mathbb{R}^{2}\!\to\!\mathbb{R}$ such that for some neighborhood of $(z^{\dagger},K_{u}^{\dagger})$, $f(z,K_{u},g(z,K_{u}))\!=\!0$. We get:
@@
$=-\frac{\alpha(z^{\dagger})^{3}\tanh^{\prime}(\alpha z^{\dagger}(u_{0}^{\dagger}+K_{u}^{\dagger}(z^{\dagger})^{2}))}{\alpha z^{\dagger}\tanh^{\prime}(\alpha z^{\dagger}(u_{0}^{\dagger}+K_{u}^{\dagger}(z^{\dagger})^{2}))}=-(z^{\dagger})^{2}.$
@@
Proposition 2 implies that one limitation of NOD is that large $K_{u}$ can make multi-stability so large and robust that solutions can get “stuck” in one of the decision attractors unless very large inputs in favor of another decision state are applied. This is illustrated in Fig. 3A, where the first (dark blue trace) and second (light blue trace) NOD differ only in their $K_{u}$ parameters $K_{u_{1}}\!>\!K_{u_{2}}\!>\!\frac{d^{3}}{3\alpha}$ but their solutions are distinctively different. At the stimulus onset ($b\!>\!0$ for $1000\!\leq\!t\!<\!$ 100), the solution of the first NOD converges to $z\!>\!0$ much more rapidly than that of the second NOD. When the input switches values ($b\!<\!0$ for $t\!\geq\!100$), the solution of the first NOD gets stuck to a positive value, whereas the solution of the second NOD is able to track the change in input sign. This example reveals a fundamental trade-off between speed/robustness (first, dark blue NOD) and flexibility (second, light blue NOD).
Instead of aiming to fine-tune the gain $K_{u}$ around some hard to define fast/robust enough yet flexible enough decision-making behavior, we use mixed-feedback principles to make the system excitable, therefore inheriting both the speed of large-$K_{u}$ NODs and the flexibility of small-$K_{u}$ NODs. The behavior of the resulting E-NOD is shown in the third line (pink trace) of Fig. 3A. By generating “decision spikes” the E-NOD is as fast as the high $K_{u}$ NOD and as flexible as the low $K_{u}$ NOD. The rest of the paper is devoted to presenting the E-NOD model, its analysis, and its multi-agent generalization.
Figure 3: Opinion solutions of NOD and E-NOD over time and associated bifurcation diagrams. (A): Trajectories of NOD (2) and E-NOD (4) for initial condition $(z,u_{0})|_{t=0}=(0.01,0.9)$ for two $K_{u}$ values. Input signal $b$ is also shown over time. (B): Bifurcation diagrams of (2) for the two values of $K_{u}$, showing the solutions in A. moving from initial state to steady state.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: The system solutions and its $(u_{s},z)$ phase portrait as the basal attention $u_{0}$ increases. For all plots, $d\!=\!1,\alpha\!=\!2,$ (thus $u_{0}^{*}=0.5$)$,K_{u}\!=\!2,K_{u_{s}}\!=\!6,\tau_{z}\!=\!2,\tau_{u_{s}}\!=\!20$. (Top): Example solutions of $u_{s}$ and $z$ over time, with initial condition as $(u_{s},z)|_{t=0}=(0.01,0.01)$ and additive Gaussian distributed white noise. (Bottom): The $u_{s}$-nullcline ($z$-nullcline) is shown as a blue (pink) line. Solid (dotted) lines indicate stable (unstable) branches of the $z$-nullcline with respect to (4a). Gray arrows denote the vector field. Black filled circles show stable equilibria, unfilled are unstable equilibria, partially filled is a saddle-node bifurcations. Crosses show saddle equilibria. Saddle-node-homoclinic cycles in (b), and limit cycles in (c) and (d) are in yellow.
III Excitable decision-making
We present the E-NOD model in Section III-A, and analyze its behavior through geometric phase-plane analysis in Section III-B.
III-A E-NOD for a Single Agent
We define E-NOD by introducing and coupling a slow regulation variable $u_{s}$ to NOD (2):
$\tau_{z}\dot{z}=-d\,z+\tanh\Big{(}(u_{0}-u_{s}+K_{u}z^{2})\cdot\left(\alpha z\right)+b\Big{)}$ (4a)
@@ III-A E-NOD for a Single Agent
where $\tau_{u_{s}}\!\gg\!\tau_{z}$ is larger by at least an order of magnitude such that $u_{s}$ increases more slowly than opinion $z$. E-NOD (4) describes dynamics with excitability: a fast positive feedback (mediated by $z$) acts to excite the system, while a slow negative feedback (mediated by $u_{s}$) regulates it back to near its ultrasensitive pitchfork singularity.
@@
The fast positive feedback in (4a) comes from the term $(u_{0}+K_{u}z^{2})(\alpha z)$, as in NOD (2). The slow negative feedback comes from the coupling of the new variable $u_{s}$. As $z$ grows in magnitude (whether positive or negative, i.e., independent of which option is preferred), $u_{s}$ also grows, although more slowly, according to (4b). When $u_{s}$ becomes large, because of the $-u_{s}(\alpha z)$ term in  (4a), it drives $z$ back towards zero. Then as $z$ decreases towards zero, $u_{s}$ also decreases back towards zero according to (4b). The result is a spike in both $z$ and $u_{s}$.
III-B Geometric Analysis of E-NOD
@@
We use phase-plane analysis to study and illustrate the spiking and decision-making behavior of E-NOD (4). To construct the phase-plane, we first compute the nullclines.
@@
The $z$-nullcline is defined as the solutions $z$ as a function of $u_{s}$ that satisfy $\dot{z}\!=\!0$ for (4a). This is the same as solving for the equilibrium solutions of (2) as a function of $u_{0}$, except now it is a function of $u_{s}$. Thus, the $z$-nullcline (pink line in Fig. 4) corresponds to the bifurcation diagram of (2), mirrored about the vertical axis and shifted right by $u_{0}$ Here, when $b\!=\!0$, the neutral solution $z\!=\!0$ is stable (unstable) for $u_{s}\!>\!u_{s}^{*}$ ($u_{s}\!<\!u_{s}^{*}$).
The $u_{s}$-nullcline (blue line in Fig. 4) is defined as the solutions that satisfy $\dot{u}_{s}\!=\!0$ in (4b), which gives the quartic function $u_{s}\!=\!K_{u_{s}}z^{4}$. This describes a quartic parabola with symmetric solutions about the $u_{s}$ axis and vertex at $u_{s}\!=\!0$. As $K_{u}$ increases, this parabola becomes narrower.
@@
We next examine the intersections of the nullclines, which determine equilibrium solutions of E-NOD (4). As illustrated in Fig. 4 the intersections depend on the value of $u_{0}$. If $b\!=\!0$, the neutral state $(u_{s},z)\!=\!(0,0)$ is always an equlibrium and two more equilibria, symmetric about $z\!=\!0$, may be present for sufficiently high $u_{0}$.
@@
Fig. 4a illustrates the phase-plane when $b\!=\!0$ and $u_{0}\!<\!u_{0}^{th}$, $u_{0}^{th}>0$ a threshold. The nullclines have only one point of intersection at the neutral state. The neutral state is stable. Trajectories will converge to and settle at this point and no excitable behavior in the decision-making will take place.
@@
Fig. 4b illustrates the phase-plane when $b\!=\!0$ and $u_{0}\!=\!u_{0}^{th}$. The nullclines shown have three points of intersection, the neutral state and two unstable equilibria symmetric about $z\!=\!0$. The neutral state is a saddle-node bifurcation with one exponentially stable eigendirection (parallel to the $u_{s}$-axis $z=0$) and one marginally unstable direction (parallel to the $z$-axis $u_{s}=0$). There are two saddle-node-homoclinic (infinite period) cycles, one diverging upward from the saddle-node and another diverging downward from it. In the absence of noise/exogenous perturbations, all trajectories asymptotically converge to the saddle-node. The presence of any, arbitrarily small, noisy perturbation, makes the trajectories escape from the saddle-node at random time instants along either the upward or downward saddle-node-homoclinic cycle, leading to large prototypical excursions in the $(u_{s},z)$ plane
@@
These large prototypical excursions are highly reminiscent of “spiking” trajectories of excitable neuronal system. By analogy, we call them “decision spikes” or “excitable decisions”. However, in contrast to neuronal spikes which happen only in one direction, decision spikes can happen in multiple directions, i.e., in as many directions as there are options. For the one-dimensional two options dynamics studied here, both upward (in favor of option 1) and downward (in favor of option 2) decision spikes are possible.
@@
When $b\!=\!0$ and $u_{0}\!>\!u^{th}$ (Fig. 4c) the three fixed point are unstable and it is possible to prove, along the same lines as [20], the existence of two limit cycles, mirror symmetric about the horizontal axis $z=0$. These limit cycles are made of repetitive decision spikes, i.e., spiking decision limit cycles. Geometric singular perturbation analysis [21, 22] provides the tools to rigorously prove the existence of these spiking decision limit cycles. Such an analysis goes beyond the scope of this paper. Instead, we leverage Fig. 4c to describe qualitatively a typical oscillatory spiking decision behavior in the presence of small noisy perturbations.
@@
Consider a trajectory with initial conditions in the rightmost part of the phase portrait. Initially, the trajectory rapidly converges to the $z=0$ axis and then slowly slides leftward approaching the neutral state. As soon as the trajectory hits this equilibrium, noisy perturbations push it either upward or downward, generating an upward or downward decision spike respectively. The decision spike trajectory brings the trajectory back to the pitchfork singularity, the next decision spike is generated, and the spiking decision cycle continues.
@@
When input $b\!\neq\!0$ and $|b|$ is sufficiently large, the $z$-nullcline unfolds accordingly to the universal unfolding of the pitchfork [17, Ch. III]. Due to the nullcline unfolding, the phase-plane geometry changes qualitatively as shown in Fig. 4d. Similar geometric singular perturbation analysis methods as those employed for the analysis of Fig. 4b and 4c reveal the existence of a unique spiking decision limit cycle associated to spiking decisions toward the option favored by the inputs (option 1, upward decision spikes in the case of Fig. 4d where $b>0$ provides evidence in favor of option 1).
@@
Observe that in the presence of informative inputs (Fig. 4d) the decision spiking frequency is higher than in the case of endogenous decision spiking oscillations (Fig. 4c). This feature is akin to spike frequency coding of input intensity in biological neural systems. In applications like robot navigation, $u_{0}$ can be controlled to avoid endogenous spiking.
IV Multi-agent excitable decision-making
It is possible to generalize the E-NOD model to the case of $N_{a}$ agents in a way analogous to how the NOD is generalized [1] but including the new slow regulation variable(s). There are two natural, non-equivalent, ways of doing this. One is to introduce a centralized regulation variable $u_{s}$, and the other is to introduce distributed regulation variables $u_{s,i}$. We focus on the distributed solution. For each $1\!\leq\!i\!\leq\!N_{a}$, agent $i$ has two state variables $z_{i}$ and $u_{s,i}$ with dynamics given by
$\textstyle\dot{z}_{i}=-d\,z_{i}+\tanh\left((u_{0}-u_{s,i}+K_{u}z_{i}^{2})\left(\sum_{k=1}^{N_{a}}a_{ik}z_{k}+b_{i}\right)\right)$ (5a)
$\dot{u}_{s,i}=-u_{s,i}+K_{u_{s}}z_{i}^{4}$ (5b)

### v3

@@ II-C Limitation on Tuning of NOD
($K_{u}$ determines region of multi-stability): Let $b\!=\!0$ and $(u_{0}^{\dagger},z^{\dagger})$ be either of the two saddle-node bifurcation points of the subcritical pitchfork of NOD (2) for $K_{u}\!>\!\frac{d^{3}}{3a}$. Then $u_{0}^{\dagger}$ is a monotonically decreasing function of $K_{u}$, i.e., $\frac{\partial u_{0}^{\dagger}}{\partial K_{u}}\!<\!0$.
@@
Let $K_{u}^{\dagger}\!>\!\frac{d^{3}}{3a}$ and $f(z,K_{u},u_{0})\!:=\!-d\,z\!+\!\tanh(az(u_{0}\!+\!K_{u}z^{2}))$. By hypothesis, $f(z^{\dagger},K_{u}^{\dagger},u_{0}^{\dagger})\!=\!0$. We have $\frac{\partial f}{\partial u_{0}}(z^{\dagger},K_{u}^{\dagger},u_{0}^{\dagger})=az^{\dagger}\tanh^{\prime}(az^{\dagger}(u_{0}^{\dagger}\!+\!K_{u}^{\dagger}(z^{\dagger})^{2}))\!\neq\!0$, since $z^{\dagger}\!\neq\!0$. Following [21], we use the implicit function theorem to show the existence of $g\!:\!\mathbb{R}^{2}\!\to\!\mathbb{R}$ such that for some neighborhood of $(z^{\dagger},K_{u}^{\dagger})$, $f(z,K_{u},g(z,K_{u}))\!=\!0$. We get
@@
$=-\frac{a(z^{\dagger})^{3}\tanh^{\prime}(az^{\dagger}(u_{0}^{\dagger}+K_{u}^{\dagger}(z^{\dagger})^{2}))}{az^{\dagger}\tanh^{\prime}(az^{\dagger}(u_{0}^{\dagger}+K_{u}^{\dagger}(z^{\dagger})^{2}))}=-(z^{\dagger})^{2}.$
@@
Figure 4: The system solutions and $(u_{s},z)$ phase portrait as the basal attention $u_{0}$ increases. For all, $d\!=\!1,\alpha\!=\!2,$ (thus $u_{0}^{*}\!=\!0.5$)$,K_{u}\!=\!2,K_{u_{s}}\!=\!6,\tau_{u_{s}}/\tau_{z}\!=\!10$. (Top): Example solutions of $u_{s}$ and $z$ over time, with initial condition as $(u_{s},z)|_{t=0}=(0.01,0.01)$ and additive Gaussian distributed white noise. (Bottom): The $u_{s}$-nullcline ($z$-nullcline) is shown as a blue (pink) line. Solid (dotted) lines indicate stable (unstable) branches of the $z$-nullcline with respect to (4a). Gray arrows denote the vector field. Black filled circles show stable equilibria, unfilled are unstable equilibria, partially filled is a saddle-node bifurcations. Crosses show saddle equilibria. Saddle-node-homoclinic cycles in (b), and limit cycles in (c) and (d) are in yellow.
Proposition 2 implies that one limitation of NOD is that large $K_{u}$ can make the region of multi-stability of NOD (2) so large and robust that solutions can get “stuck” in one of the decision attractors unless very large inputs in favor of another decision state are applied. This is illustrated in Fig. 3A, where the first (dark blue) and second (light blue) NOD differ only in their $K_{u}$ parameters ($K_{u_{1}}\!>\!K_{u_{2}}\!>\!\frac{d^{3}}{3a}$) but their solutions are distinctively different. At the stimulus onset ($b\!>\!0$ for $1000\!\leq\!t\!<\!$ 100), the solution of the first NOD converges to $z\!>\!0$ much more rapidly than that of the second NOD. However, when the input switches values ($b\!<\!0$ for $t\!\geq\!100$), the solution of the first NOD gets stuck at a positive value, whereas the solution of the second NOD is able to track the change in input sign. This example reveals a fundamental trade-off between speed/robustness (first, dark blue NOD) and flexibility (second, light blue NOD).
Instead of aiming to fine-tune the gain $K_{u}$ around a hard-to-define fast/robust enough yet flexible enough decision-making behavior, we use mixed-feedback principles to make the system excitable, inheriting both the speed of large-$K_{u}$ NODs and the flexibility of small-$K_{u}$ NODs and imparting system agility. The behavior of the resulting S-NOD is shown in pink in Fig. 3A. By generating “decision spikes” the S-NOD is as fast as the high $K_{u}$ NOD and as flexible as the low $K_{u}$ NOD. In what follows, we present the S-NOD model, its analysis, and its multi-agent generalization.
III Agile Decision-Making: S-NOD
We present and analyze the Spiking Nonlinear Opinion Dynamics (S-NOD) model.
III-A S-NOD for a Single Agent and Two Options
We define S-NOD by introducing a slow regulation variable $u_{s}$ to NOD (2), as in the fast-positive, slow-negative mixed-feedback structure of excitable systems [1, 2, 3, 4]:
$\tau_{z}\dot{z}=-d\,z+\tanh\Big{(}(u_{0}-u_{s}+K_{u}z^{2})\cdot\left(az\right)+b\Big{)}$ (4a)
@@ III-A S-NOD for a Single Agent and Two Options
where $\tau_{u_{s}}\!\gg\!\tau_{z}$ is larger by at least an order of magnitude such that $u_{s}$ responds more slowly than $z$. S-NOD as in (4) describes dynamics with excitability: a fast positive feedback (mediated by $z$) acts to excite the system, while a slow negative feedback (mediated by $u_{s}$) regulates it back to near the ultrasensitive pitchfork singularity (as seen in Fig. 2).
@@
The fast positive feedback in (4a) comes from the term $(u_{0}+K_{u}z^{2})(az)$, as in NOD (2). The slow negative feedback comes from the coupling of the new variable $u_{s}$. As $z$ grows in magnitude (whether positive or negative, i.e., independent of which option is preferred), $u_{s}$ also grows, although more slowly, according to (4b). When $u_{s}$ becomes large, it drives $z$ back towards zero because of the $-u_{s}(az)$ term in  (4a). As $z$ decreases towards zero, $u_{s}$ also decreases back towards zero according to (4b). The result is a spike in $z$ and $u_{s}$. This is a limit cycle, which bounds $z$ and $u_{s}$, and is analyzed next.
The S-NOD equations (4) can be generalized to $N_{o}\geq 2$ options analogously to the generalization of NOD to $N_{o}$ options [11, 16]. We will investigate this in future work.
III-B Geometric Analysis of Single-Agent, Two-Option S-NOD
@@
We use phase-plane analysis to study and illustrate the spiking and decision-making behavior of S-NOD (4). To construct the phase-plane, we first compute the nullclines.
@@
The $z$-nullcline is defined as the solution pairs $(u_{s},z)$ that satisfy $\dot{z}\!=\!0$ for (4a). This is equivalent to solving for the equilibrium solutions of (2) as a function of $u_{0}$ as in Section II-B. Thus, the $z$-nullcline (pink in Fig. 4) is analogous to the bifurcation diagram of (2), mirrored about the vertical axis and shifted right by $u_{0}$, with $u_{s}^{*}\!=\!u_{0}\!-\!d/a\!=\!u_{0}\!-\!u_{0}^{*}$. When $b\!=\!0$, the neutral solution $z\!=\!0$ is a stable (unstable) equilibrium of (4a) for $u_{s}\!>\!u_{s}^{*}$ ($u_{s}\!<\!u_{s}^{*}$). The $u_{s}$-nullcline (blue in Fig. 4) is defined as the solution pairs $(u_{s},z)$ that satisfy $\dot{u}_{s}\!=\!0$ in (4b), which gives the quartic parabola $u_{s}\!=\!K_{u_{s}}z^{4}$. The larger $K_{u}$, the more narrow this parabola is.
@@
The intersections of the nullclines determine equilibrium solutions of S-NOD (4) and, as shown in Fig. 4, depend on the value of $u_{0}$. If $b\!=\!0$, the neutral solution $(u_{s},z)\!=\!(0,0)$ is always an equilibrium. Two more equilibria, symmetric about $z\!=\!0$, may be present for high enough $u_{0}$.
@@
Fig. 4a depicts the phase-plane when $b\!=\!0$ and $u_{0}\!<\!u_{0}^{*}$. The nullclines have one point of intersection at the neutral solution. The neutral solution is stable. Trajectories will converge to and settle at this point and no excitable behavior in the decision-making will take place.
@@
Fig. 4b depicts the phase-plane when $b\!=\!0$ and $u_{0}\!=\!u_{0}^{*}$. The nullclines have three points of intersection, the neutral solution and two unstable equilibria symmetric about $z\!=\!0$. The neutral solution is a saddle-node bifurcation with one exponentially stable eigendirection (along $z\!=\!0$) and one marginally unstable eigendirection (along $u_{s}\!=\!0$). There are two saddle-node-homoclinic (infinite period) cycles, diverging either upward and downward from the saddle-node. In the absence of noise/exogenous perturbations, all trajectories asymptotically converge to the neutral solution. The presence of any arbitrarily small noise makes the trajectory escape from the neutral solution at random time instants along either the upward or downward saddle-node-homoclinic cycle, leading to large prototypical excursions in the $(u_{s},z)$ plane.
@@
These large prototypical excursions resemble of “spiking” trajectories of excitable neuronal system. By analogy, we call them “decision spikes” or “excitable decisions”. In contrast to neuronal spikes which happen in only one direction, decision spikes can happen in as many directions as there are options. For the one-dimensional two-option dynamics studied here, both upward (in favor of option 1) and downward (in favor of option 2) decision spikes are possible.
@@
Fig. 4c depicts the phase-plane when $b\!=\!0$ and $u_{0}\!>\!u_{0}^{*}$. Three fixed points are unstable and it is possible to prove, along the same lines as [22], the existence of two limit cycles, symmetric about the horizontal axis $z=0$. These limit cycles are made of repetitive decision spikes, i.e., spiking decision limit cycles. Geometric singular perturbation analysis [23, 24] provides the tools to rigorously prove the existence of these spiking decision limit cycles. Such an analysis goes beyond the scope of this paper. Instead, we leverage Fig. 4c to describe qualitatively a typical oscillatory spiking decision behavior in the presence of small noisy perturbations.
@@
Consider Fig. 4c and a trajectory with a large initial condition in $u_{s}$. Initially, the trajectory rapidly converges to the $z=0$ axis, then slowly slides leftward approaching the neutral solution. As soon as the trajectory hits this equilibrium, noisy perturbations push it either upward or downward, generating an upward or downward decision spike respectively. The decision spike trajectory brings the trajectory back to the pitchfork singularity, the next decision spike is generated, and the spiking decision cycle continues.
@@
When input $b\!\neq\!0$ and $|b|$ is sufficiently large, the $z$-nullcline unfolds accordingly to the universal unfolding of the pitchfork [19, Ch. III]. Due to the nullcline unfolding, the phase-plane geometry changes qualitatively as shown in Fig. 4d. Similar geometric singular perturbation analysis methods as those employed for the analysis of Fig. 4b and 4c reveal the existence of a unique spiking decision limit cycle associated to spiking decisions toward the option favored by the inputs (e.g., upward decision spikes in the case of Fig. 4d where $b>0$ provides evidence in favor of option 1).
@@
Observe that in the presence of informative inputs (Fig. 4d), the decision spiking frequency is higher than in the case of endogenous decision spiking oscillations (Fig. 4c). This feature is similar to spike frequency indicating input intensity in neural systems. In applications like robot navigation, $u_{0}$ can be controlled to avoid endogenous spiking.
IV Agile Multi-Agent Decision-Making: S-NOD
IV-A S-NOD for Multiple Agents and Two Options
Figure 5: Response of three agents with the same communication network but for opposite signs on graph edge weights and an input $b_{1}$ applied only to agent $1$. For $i\neq k$, (a) $a_{ik}=+0.1$ and (b) $a_{ik}=-0.1$. Parameters are $a_{ii}=1$, $d=1$, $K_{u}=2.3$, $K_{u_{s}}=16$, $u_{0}=0.9$, $\tau_{u_{s}}/\tau_{z}=20$.
We can generalize the single-agent S-NOD equations (4) to the case of $N_{a}$ agents in the same way that NOD generalizes to $N_{a}$ agents [11]. The multi-agent NOD models the decision-making process of multiple agents sharing and influencing one another’s opinions over a communication network. Examples include agents choosing how they distribute their time over two resource patches or deciding whether to move to the left or right when navigating a cluttered space, all while integrating information from other agents’ opinions. In the multi-agent S-NOD, each agent $i$ has two state variables $z_{i}$ and $u_{s,i}$ with dynamics given by
$\tau_{z}\dot{z}_{i}\!=\!-d\,z_{i}\!+\!\tanh\!\left(\!(\!u_{0}\!-\!u_{s,i}\!+\!K_{u}z_{i}^{2})\!\left(\!\sum_{k=1}^{N_{a}}\!a_{ik}z_{k}\!\right)+\!b_{i}\!\right)$ (5a)
$\tau_{u_{s}}\dot{u}_{s,i}=-u_{s,i}+K_{u_{s}}z_{i}^{4}$ (5b)